OpenStudy (anonymous):

for 0 less than or equal to x less than pi/2, an antiderivative of 2tanx is?

OpenStudy (anonymous):

Do you have to derive it yourself?

OpenStudy (anonymous):

I have to find the anti derivative of it. My friend told me that the answer if ln(sec2x) but I'm not sure how to get that

OpenStudy (freckles):

have you guys done solving integrals by substitution ?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

so you know you can write tan(x) as sin(x)/cos(x) and sub u=cos(x)

OpenStudy (freckles):

\[\int\limits \frac{\sin(x)}{\cos(x)}dx \\ u=\cos(x) \\ \frac{du}{dx}=-\sin(x) \\ -du= \sin(x) dx\]